OpenStudy (anonymous):

i need to make up and equation like this... For example, if John takes 2 hours to mow his lawn and it takes his sister, Maria, 4 hours to mow the same lawn, how long would it take John and Maria to mow the lawn together?... And then be able to write a rational equation based upon it.

OpenStudy (anonymous):

Didn't I already answer this? Oh, well I guess I can do it again: Cato fills a swimming pool in 8 hours and it takes his aunt Alma 10 hours to do the same. If they fill it together, how long will it take?

OpenStudy (anonymous):

waiittt someone else might have asked for this...please..dont give me a example u already used bc i will get into troubled they will think its palgierized

OpenStudy (anonymous):

No, this example I've never given out before. Don't worry.

OpenStudy (anonymous):

okay! ANd i will try and solve hold on

OpenStudy (anonymous):

so its going to be....1/8+1/10=1/x....?

OpenStudy (anonymous):

No, x and y is the two different times used to carry out the task. In this case 8 and 10.

OpenStudy (anonymous):

One of these: x= 8 and y= 10 or x= 10 and y= 8

OpenStudy (anonymous):

wait im confused so i wrote the equation wrong?

OpenStudy (anonymous):

are you doing online school? flvs?

OpenStudy (anonymous):

yes and i hate it .... !...it was a class from last year and its holding me back from graduating

OpenStudy (anonymous):

Sorry, I wrote it wrong too: \[(x*y)/(x+y)\] Where x and y is the two different times used to preform a task.

OpenStudy (anonymous):

:/ yea im doing the same ! hey i just turn in that work so how about ill send u mine and just change the names and maybe the numbers so it wont be the same cuz we probably have the same teacher

OpenStudy (anonymous):

this is what i sent : Suncoast Community High School is having a car wash fundraiser for their senior class trip. It takes Joanne four hours to wash a car, while Maria can wash a car in three hours. If they work together, how much time would it take them to wash and wax one car? In one hour, Joanne can wash ¼ of a car, and if we represent time with an x variable, Joanne can wash x/4 cars. Same for Maria x/3 represents how many cars Maria can wash in x hours. Also since both Maria and Joanne are working together we add the two expressions. Furthermore, we are trying to find out how many hours it takes them to wash and wax one car. So, the rational expression will equal 1 and we solve for x. so, the rational expression for this work problem is x/4+x/3=1 to solve: first, find the least common denominator, in this case is 12. to make all the denominators 12, we multiply x(3)/4(3)+x(4)/3(4)=1(12), so were left with 3x+4x=12, and since the denominators are all 12, we just eliminate them and solve for x. So, 7x=12, divined 12/7 and x=12/7. It will take Joanne and Maria 12/7, or more accurately, 1.7 hours to wash and wax one car if they work together.

OpenStudy (anonymous):

hiope it helps good luck :)

OpenStudy (anonymous):

suncoast? lol....im gussin we r not far from eachother i go to pbc

OpenStudy (anonymous):

And no thanks he can just work this out with me but that was sweet of u!

OpenStudy (anonymous):

@prebz (8x10)(8+10)??

OpenStudy (anonymous):

i used to go there last year... now im doing online .. oh ur welcome :)

OpenStudy (anonymous):

Yea, but with a dividing line between the two parentheses

OpenStudy (anonymous):

(8*10):(8+10)

OpenStudy (anonymous):

OO OKAY!

OpenStudy (anonymous):

so thats the equation?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay..now let me see and solve

OpenStudy (anonymous):

4.4?

OpenStudy (anonymous):

Yup, or if you wanna be iffy about it: 4 hours 26 minutes 40 seconds :)

OpenStudy (anonymous):

haha thanks

OpenStudy (anonymous):

can u solve this.....1/2+1/4=1/x

OpenStudy (anonymous):

Since 1/2=2/4 We get: 2/4+1/4=1/x 3/4=1/x x*(3/4)=1 x=(1/1):(3/4) x=4/3 or 1.333...

OpenStudy (anonymous):

thanksss u just made my whole night so muchh easier on me

OpenStudy (anonymous):

No problem